Jenny James is paid a weekly salary of $250 plus a 4% commission on sales. Last...

  1. Jenny James is paid a weekly salary of $250 plus a 4% commission on sales. Last week her sales were $10,876. What were her earnings for the week? (2 marks)
  1. Calculate the total commission paid to each salesperson. (4 marks)

Name

Total Sales

Commission

Total Commission

5% on first $5000

7.5% on

Next $25,000

13% on all Additional sales

Ross, J

$195,650

Vilia, K

$237,500

  1. Mary Eisen worked 7 hours a day from Monday to Thursday at a regular pay rate of $11.25 per hour. Friday was a statuary holiday and she worked 10 hours at double her regular hourly pay. What was Mary’s gross pay for the week? (2 marks)
  1. Calculate the net pay for each employee. (Overtime is at ‘time and a half’) (4 marks)

           

Emp. No.

Hours

Hourly Rate

Earnings

Deductions

Net Pay

Reg.

O.T

Reg.

O.T.

Gross

Income Tax

E.I.

C.P.P.

Total

13

36

$9.50

-

44.25

8.72

20.90

14

40

2

$15.45

127.25

16.94

20.90

  1. The following workers are paid $2.25 for a lathe-turned spindle. A quota of 200 spindles per week is set. In any given week any spindles turned in excess of 200 are paid at a rate of $3.01 per spindle. Calculate the number of spindles turned and the gross pay for each worker. (3 marks)

Name

Number of Pieces

Total Pieces

Gross Pay

Mon.

Tues.

Wed.

Thur.

Fri

Atkins, J

40

38

39

40

41

Degrande, M

40

41

47

48

46

please show the solution...
Answer #1

1) Jenny James:

Fixed Weekly salary = $ 250

Commission = 4% of sales = 4/100*10876 = $435.04

Total earnings for the week = Fixed salary + commission = 250 + 435.04 = $ 685.04

2) Commission paid to each salesperson:

a) Ross J

Sales = $1,95,650

Total commission of Ross J = (5% of 5000) + (7.5% of 25,000) + (13% of 1,65,650) = $23,659.5

b) Vilia K

Sales = $ 2,37,500

Total commission of Ross J = (5% of 5000) + (7.5% of 25,000) + (13% of 2,07,500) = $29,100

3) Mary Eisen

Mary's total pay from Monday to Thursday = 4 * 7 * 11.25 = $315

Mary's total pay from Monday to Thursday = 10 * 2 * 11.25 = $225

Mary's gross pay for the week = 315 + 225 = $ 540

4) Table alignment is not clear

5) No. of spindles turned and gross pay for each worker

a) Mon Atkins

Total spindles turned for the week = 40+38+39+40+41 = 198

Gross Pay = 198 * 2.25 = $445.5

b) Degrande M

Total spindles turned for the week = 40+41+47+48+46 = 222

Gross Pay = (200 * 2.25) + (22*3.01) = $516.22